# Speeding up quantum measurement using space-time trade-off

### v1

@@ V Conclusion
In this paper we have presented a general scheme for speeding up quantum measurement while maintaining quality. This is achieved by means of a space-time trade-off, using additional ancillary qubits in place of longer measurement time. We presented the main idea using ideal Poisson measurement statistics but emphasise that the scheme is much more general than this. In particular, by using characteristic noise models for the CNOTs and decoherence, the scheme was shown to be robust to imperfections while simultaneously able to achieve a “super-improvement” in speed-up, since it acts to suppresses decoherence, allowing for a higher achievable maximum quality. While we have illustrated the scheme with systems of qubits, it should be clear from the discussions above that the scheme is equally applicable to higher dimensional systems.
@@ VII Acknowledgements
VIII Appendices
VIII.1 Appendix on noise model
@@
For the CNOT noise, we consider a simple model inspired by neutral atom quantum computers. In this model, whenever the system is in state |$\ket{1}$ and a CNOT is applied, there is a chance of failure, which causes the system to decay to |$\ket{0}$. More formally, we define a probability $p$, that immediately before a CNOT is implemented, |$\ket{1}$ will be sent to |$\ket{0}$. If this happens, the CNOT can not flip the state |$\ket{0}$ back to |$\ket{1}$. The noisy CNOT channel $\widetilde{C_{p}}$ acts on the basis states, where the first qubit is the control qubit, as
@@
Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{0}$ and |$\ket{1}$ of the system qubit. Furthermore, we assume our ancillary qubits are initialised perfectly in |$\ket{0}$ so we are only interested in how the noisy CNOT gates will act on the states |$\ket{00}$ and |$\ket{10}$. Thus (VIII.1) defines the action on all states that occur in the scheme.
When the system is in the state |$\ket{0}$, and we perform $N$ measurements, the initial state will be |$\ket{0}^{\otimes N}$. In this case, the action of $\widetilde{C_{p}}$ is to leave the state of all $N$ systems as |$\ket{0}^{\otimes N}$. When the system is in the state |$\ket{1}$, the initial state is |$\ket{1}\ket{0}^{\otimes(N-1)}$. Our scheme only depends upon the total number of photons emitted from each qubit (and does not depend upon which qubit emitted which photon). We thus only need to calculate the probability distribution for the total number of qubits in state |$\ket{1}$ after applying the $\widetilde{C_{p}}$ gates.
There are numerous ways the gates can be compiled and different compilations will give different probability distributions for the number of qubits in state |$\ket{1}$. These different compilations are explored in more detail in the supplementary material. Let us denote by $\mathcal{T}^{(N)}_{p}(q)$ the probability distribution of the random variable $Q^{(N)}_{p}$ defining the number of qubits in state |$\ket{1}$ after applying the noisy CNOT gates. Any qubits not in state |$\ket{1}$ are in state |$\ket{0}$. Combining this with the probability distribution for $N$ qubits emitting $k$ photons (7), we find the resultant distribution using our scheme with noisy CNOT gates to be
$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k)$
@@
For the decoherence noise, we consider a model whereby qubits in state |$\ket{1}$ probabilistically decay at a constant rate $\lambda$ to state |$\ket{0}$. After this happens, the qubit will emit photons at a different rate so the resultant distribution of photons will not be the same as a perfect Poisson. We are interested in the probability distribution for the number of photons emitted $k$ in time $t$. We model this behaviour by finding the probability a qubit in state |$\ket{1}$ will emit $k^{\prime}$ photons in time $t^{\prime}$ before decaying to |$\ket{0}$ and then emitting $k-k^{\prime}$ photons in time $t-t^{\prime}$, then integrating over $t^{\prime}$ and summing over $k^{\prime}$. We find the resultant probability distribution to be
@@
Figure 3: Increase in peak SNR from Fig. 2 (c) and (d). These Plots are generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1) with the cascade compilation of noisy CNOT gates defined in the Supplementary material. We use input parameters $\mu_{0}=3.5$, $\mu_{1}=14$, $\lambda=0.0041$ and $p=0.01$. We plot the peak SNR $S_{max}$ for $N=1$ to $N=10$ qubits. While more qubits result in a higher peak SNR, this trend flattens out for higher qubit numbers.
Figure 4: Speed up from Fig. 2 (c) and (d). Plot generated using the complete noise model of the Poisson distributed measurement statistics (VIII.1) with the cascade compilation of noisy CNOT gates defined in the Supplementary material. We use input parameters $\mu_{0}=3.5$, $\mu_{1}=14$ and $\lambda=0.0041$. We plot the ratio $t^{(1)}_{S}/t^{(N)}_{S}$ for two different values of $p$, at $S=8$, for $N=1$ to $N=10$ qubits and the linear relationship $t^{(1)}_{S}/t^{(N)}_{S}=N$ (lime dashed line) achieved by the perfect Poisson distribution (8). The value $p=0.01$ (blue dots) is used in Fig. 2 (c) and (d). The points are increasing, which shows we get a speed up, however they are below the dashed line so the speed-up is worse than the linear case. On the other hand, the values for $p=0.001$ (red crosses) are above the dashed line which demonstrates that, for CNOTs with a low enough failure rate, we can beat the decoherence of the qubit; it is possible to achieve a “super-improvement”.
Our complete noise model consists of combining the noisy CNOT protocol (VIII.1) with the decaying Poisson distribution (15). Altogether, we we thus find that the probability distributions for the number of photons, using our scheme with $N$ systems measured, is given by
$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k)$
@@
These are used to produce the results in Fig. 2 (c) and (d).

### v2

@@ VI Conclusion
In this paper we have presented a general scheme for speeding up quantum measurement while maintaining quality. This is achieved by means of a space-time trade-off, using additional ancillary qubits in place of longer measurement time. We presented the main idea using ideal Poisson measurement statistics but emphasise that the scheme is much more general than this. Moreover, by using characteristic noise models for the CNOTs and decoherence, the scheme was shown to be robust to imperfections while simultaneously able to achieve a “super-improvement” in speed-up, since it acts to suppresses decoherence, allowing for a higher achievable maximum quality. While we have illustrated the scheme with systems of qubits, it should be clear from the discussions above that the scheme is equally applicable to higher dimensional systems.
Our case study was framed around state dependent fluorescence qubit measurements which are particularly appropriate to trapped-ion and cold-atom qubits. But, we would like to reiterate that our ideas are much more general. For example, our ideas also apply, when using multiple detectors, to superconducting qubits. We give a very brief discussion of this in the Supplementary Material. We also believe, just as in the earlier discussions, that we can embody our scheme for superconductors using a single detector; this is the subject of future work.
@@ VIII Acknowledgements
IX Appendix on noise model
@@
(i) For the CNOT noise, we consider a simple model inspired by neutral-atom quantum computers. In this model, whenever the system is in state |$\ket{1}$ and a CNOT is applied, there is a chance of failure, which causes the system to decay to |$\ket{0}$. More formally, we define a probability $p$, that immediately before a CNOT is implemented, |$\ket{1}$ will be sent to |$\ket{0}$. If this happens, the CNOT can not flip |$\ket{0}$ back to |$\ket{1}$. The noisy CNOT channel $\widetilde{C_{p}}$ acts on the basis states, where the first qubit is the control qubit, as
@@
Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{\psi}=\ket{0}$ and |$\ket{\psi}=\ket{1}$ of the system qubit. Furthermore, we assume our ancillary qubits are initialised perfectly in |$\ket{0}$ so we are only interested in how the noisy CNOT gates will act on the states |$\ket{00}$ and |$\ket{10}$. Thus (IX) defines the action on all states that occur in the scheme.
When the system is in the state |$\ket{\psi}=\ket{0}$, and we have $N-1$ ancillary qubits, the initial state will be |$\ket{0}^{\otimes N}$. In this case, the action of $\widetilde{C_{p}}$ is to leave the state of all $N$ systems as |$\ket{0}^{\otimes N}$. When the system is in the state |$\ket{\psi}=\ket{1}$, the initial state is |$\ket{1}\ket{0}^{\otimes(N-1)}$. Since our scheme only depends upon the total number of photons emitted from each qubit (and does not depend upon which qubit emitted which photon), we only need to calculate the probability distribution for the total number of qubits in state |$\ket{1}$ after applying the $\widetilde{C_{p}}$ gates.
Figure 3: Increase in peak SNR from Fig. 2c and 2d. This figure is extra detail for Figs. 2c and 2d and uses the same parameters. We plot the peak SNR $S_{max}$ for $N=1$ to $N=10$ qubits. While more qubits results in a higher peak SNR, this trend flattens out for higher qubit numbers before eventually dropping. The slight unevenness in the plot is an odd/even effect of the cascade CNOT compilation, see Supplementary Material.
Figure 4: Speed up from Fig. 2c and 2d. This figure is extra detail for Figs. 2c and 2d and uses the same parameters except for $p$, for which we plot two values. We plot the ratio $t^{(1)}_{S}/t^{(N)}_{S}$ for two different values of $p$, at $S=8$, for $N=1$ to $N=10$ qubits and the linear relationship $t^{(1)}_{S}/t^{(N)}_{S}=N$ (lime dashed line) achieved by the perfect Poisson distribution (6). The value $p=0.01$ (blue dots) is used in Figs. 2c and 2d. The points are increasing, which shows we get a speed up, however they are below the dashed line so the speed-up is worse than the linear case. On the other hand, the values for $p=0.001$ (red crosses) are above the dashed line which demonstrates that, for CNOTs with a low enough failure rate, we can beat the decoherence of the qubit; it is possible to achieve a “super-improvement”.
There are numerous ways to compile the gates and different compilations give different probability distributions for the number of qubits in state |$\ket{1}$, given an input state |$\ket{\psi}=\ket{1}$. These compilations are explored in more detail in the Supplementary Material. Let us denote by $\mathcal{T}^{(N)}_{p}(q)$ the probability distribution of the random variable $Q^{(N)}_{p}$ defining the number of qubits in state |$\ket{1}$ after applying the noisy CNOT gates. Any qubits not in state |$\ket{1}$ are in state |$\ket{0}$. Combining this with the probability distribution for $N$ qubits emitting $k$ photons (5), we find the resultant distribution using our scheme with noisy CNOT gates to be
$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k),$
@@
(ii) For the decoherence, we consider a model whereby qubits in state |$\ket{1}$ probabilistically decay at a constant rate $\lambda$ to state |$\ket{0}$. After this happens, the qubit will emit photons at a different rate so the resultant distribution of photons will not be the same as a perfect Poisson. We are interested in the probability distribution for the number of photons emitted $k$ in time $t$. We model this behaviour by finding the probability that a qubit in state |$\ket{1}$ will emit $k^{\prime}$ photons in time $t^{\prime}$ before decaying to |$\ket{0}$ and then emitting $k-k^{\prime}$ photons in time $t-t^{\prime}$, then integrating over $t^{\prime}$ and summing over $k^{\prime}$. We find the resultant probability distribution to be
@@
Our complete noise and decoherence model consists of combining the noisy CNOT protocol (IX) with the decaying Poisson distribution (12). Altogether, we find that the probability distributions for the number of photons emitted, using an $N$ qubit scheme, is given by
$\mathcal{P}_{\ket{0},t}^{(N)}(k)$ $=L_{N\mu_{0}t}(k),$
@@
These are used to produce the results in Figs. 2c and 2d where we use the cascade CNOT compilation, found in the Supplementary Material, for $\mathcal{T}^{(N)}_{p}(q)$.